Type The following information pertains to Stark Corporation Month Jan Feb Mar Apr May June sales...

Type

The following information pertains to Stark Corporation

Month

Jan

Feb

Mar

Apr

May

June

sales and Purchases

$30,000    $10000

34,000 12000

38,000 14000

42,000    16000

48,000    18000

60,000 20000

Cash is collected from customers in the following manner

30% Month of sale*

50% Month following sale

15% Two months following sale

5% Amount uncollectible

*Customers paying in the month of sale receive a 2% cash discount.

40% of purchases are paid for in cash in the month of purchase, and the balance is paid the

following month.

Prepare a summary of cash collections for the 2nd quarter.
Answer #1

Please find below table useful to compute desired results: -

А B 2 April 3 February =34000*15% 4 March =38000*50% 5 April =42000*30%-42000*30%*2% 6 May 7 June 8 Total =SUM(B3:B7) Summary

End results would be as follows: -

A B C D Summary of Cash Collections for the 2nd Quarter 2 April May June Total Cash Collection 3 February 5.100 5.100 4 March